# PV-S3: Advancing Automatic Photovoltaic Defect Detection using Semi-Supervised Semantic Segmentation of Electroluminescence Images

### v1

@@ 5 Conclusion, Limitations and Future Work
In conclusion, our research introduces PV-S3, a semi-supervised learning framework for defect detection in PV modules. PV-S3 effectively utilizes both labeled and unlabeled data, achieving accurate detection despite limited labeled samples. The proposed approach demonstrates significant potential in automating defect detection, reducing manual efforts, and enhancing the reliability of solar PV systems. The proposed approach provides mean IoU of 67%, precision of 88.75%, recall of 72.50%, and F1-score of 78.50% on one of the largest defect detection dataset with merely 20% labeled samples, significantly outperforming existing state-of-the-art method which uses 100% labeled samples.
However, our research also highlights certain limitations such as difficulty in boundary detection. While PV-S3 provides effective guidance for defect classification, accurately delineating defect boundaries can be challenging due to the defect’s complex and irregular nature.
Looking ahead, future research should focus on enhancing detection accuracy and robustness which could lead to development of tools as used for other energy tasks [51, 10]. Additionally, efforts to collect diverse unlabeled data and employ active learning strategies could enhance the model’s generalization capabilities. Furthermore, extending our approach to encompass a broader range of defect types in solar PV modules would render the system more comprehensive and adaptable to real-world scenarios. Such advancements could significantly contribute to the automation of defect detection, reducing manual labor, and ultimately enhancing the reliability of solar PV systems.

### v3

@@ 5 Conclusion, Limitations and Future Work
In conclusion, our research introduces PV-S3, a semi-supervised learning framework for defect detection in PV modules. PV-S3 effectively utilizes both labeled and unlabeled data, achieving accurate detection despite limited labeled samples. The proposed approach demonstrates significant potential in automating defect detection, reducing manual efforts, and enhancing the reliability of PV systems. The proposed approach provides mean IoU of 67%, precision of 72.50%, recall of 78.90%, and F1-score of 78.15% on one of the largest defect detection dataset with merely 20% labeled data, significantly outperforming existing state-of-the-art methods which use 100% labeled data.
Our research also highlights certain limitations such as difficulty in boundary detection. While PV-S3 provides effective guidance for defect classification, accurately delineating defect boundaries can be challenging due to the defect’s complex and irregular nature.
Looking ahead, future research should focus on enhancing detection accuracy and robustness which could lead to development of tools as used for other energy tasks [46, 9]. Additionally, efforts to collect diverse unlabeled data and employ active learning strategies could enhance the model’s generalization capabilities. Furthermore, extending our approach to encompass a broader range of defect types in PV modules would render the system more comprehensive and adaptable to real-world scenarios. Such advancements could significantly contribute to the automation of defect detection, reducing manual labor, and ultimately enhancing the reliability of solar PV systems.
@@
Data Availability
The datasets used in this study are publicly available. The UCF-EL dataset can be accessed at https://github.com/ucf-photovoltaics/UCF-EL-Defect and the CSB dataset can be accessed at https://github.com/TheMakiran/BenchmarkELimages. These datasets are freely available for research purposes.